# Relaxed Recursive Transformers: Effective Parameter Sharing with Layer-wise LoRA

### v1

@@ 4 Related Work
Cross-layer parameter sharing has proven to be an effective method for achieving parameter efficiency in deep learning models such as RNNs (Sherstinsky, 2018; Graves, 2016b), CNNs (Eigen et al., 2014; Savarese and Maire, 2019; Guo et al., 2019), and the popular Transformer architecture. The Universal Transformer (Dehghani et al., 2019), a recurrent self-attentive model, demonstrated superior performance to non-recursive counterparts with significantly fewer parameters. This cross-layer parameter sharing approach has subsequently been explored in various tasks, including language understanding  (Lan et al., 2020), language modeling (Bai et al., 2019), and machine translation, through stacking a single layer (Dabre and Fujita, 2019), tying encoder (Milbauer et al., 2023) and decoder components (Xia et al., 2019), or partially tying layers (Takase and Kiyono, 2023). These methods often claim to achieve comparable performance with more compact models and increased computational speed, while also setting the ground for effective adaptive compute solutions (Dehghani et al., 2019; Graves, 2016b; Schuster et al., 2021).
@@
In this paper, we also discuss how Recursive Transformers can be well suited for early-exiting techniques to accelerate decoding in LLMs. The inherent recursive structure readily enables early-exiting for individual responses within a large serving batch, which is often a practical limitation of such techniques. Vanilla Transformers encounter a synchronization issue with early-exiting, where the model must forward all layers if even a single token in a batch requires full processing (exited tokens must wait for them). Several approaches attempt to exploit this idle time by computing missing KV caches for exited tokens in later layers, which are essential for subsequent sequence generation. These techniques include state propagation (Schuster et al., 2022; Elbayad et al., 2020), SkipDecode (Del Corro et al., 2023), and parallel decoding (which can be combined with Speculative Decoding) (Bae et al., 2023; Elhoushi et al., 2024; Liu et al., 2024; Chen et al., 2024b; Tang et al., 2024). Nevertheless, the heterogeneous parameters across varying model depths still hinder the efficient progression of exited tokens to subsequent sequences. In contrast, our Recursive Transformers enable parallel computation for tokens at different depths and sequences (in a continuous depth-wise batching paradigm)—also allow for parallel computation of missing KV caches with minimal overhead during the memory-bounded decoding phase.
@@ Appendix B Parameter Sharing Strategy
Takase and Kiyono (2023) discuss three strategies for partial layer tying in Transformer models, as depicted in Figure B.1. The SEQUENE strategy is the simplest, assigning the same parameters to consecutive layers. The CYCLE strategy repeatedly stacks a single block of unique layers to achieve the desired depth. Meanwhile, the CYCLE (REV) strategy stacks the lower layers in reverse order for the remaining layers. We specifically utilized the CYCLE strategy due to its superior compatibility with early-exiting, which can amplify the throughput of Recursive Transformers through continuous depth-wise batching.
@@ Appendix F Expanded Results of Initialization Methods for Looped Layers
The Average initialization method exhibited notably poor performance, particularly when applied to the Gemma model. We hypothesized that this could be attributed to instability in the model’s learned distribution, potentially arising from averaging of normalization layer weights. To investigate this further, we experimented with three different methods for initializing normalization weights, as outlined in Figure 1(b): averaging weights (Norm-avg), selecting weights from a single layer (Norm-choice), and zero initialization (Norm-zero). The performance trend observed among these methods varied across different model architectures. However, zero initialization of normalization layers resulted in a huge performance drop in certain architectures like TinyLlama and Pythia. Conversely, we observed no big difference between averaging and single-layer selection, suggesting that any form of distillation of the normalization weights appears to be sufficient for maintaining performance.
@@ Appendix G Expanded Results of Relaxed Recursive Transformers
Appendix H Relaxation of Parameter Sharing with Prefix Tuning
As discussed in §2.3, we employed LoRA modules to relax parameter sharing. However, efficiently computing multiple LoRA modules requires specialized kernels and sequential computations of the base layers and LoRA modules, incurring additional computational costs. Consequently, we explored replacing layer-specific prompts (Liu et al., 2021) as an alternative. In this approach, prompts specific to each layer are integrated as prefix tokens, being utilized to generate layer-wise key and value states for self-attention computation. This approach is significantly more amenable to parallel computation, leading to reduced computational overhead.
Table H.1 summarizes performance of the prefix tuning method. Due to the reliance on small learnable prompts, performance gains were limited. Additionally, without a mechanism to leverage the original pretrained weights, the performance of prefix tuning was significantly lower (52.1% vs. 47.6% with the Average method in 1.07B model size). While offering parallel computation advantages, further research is needed to enhance its effectiveness.

### v2

@@ 4 Related Work
Cross-layer parameter sharing has proven to be an effective method for achieving parameter efficiency in deep learning models such as RNNs (Sherstinsky, 2018; Graves, 2016b), CNNs (Eigen et al., 2014; Savarese and Maire, 2019; Guo et al., 2019; Shen et al., 2022), and the popular Transformer architecture. The Universal Transformer (Dehghani et al., 2019), a recurrent self-attentive model, demonstrated superior performance to non-recursive counterparts with significantly fewer parameters. This cross-layer parameter sharing approach has subsequently been explored in various tasks, including language understanding  (Lan et al., 2020), language modeling (Bai et al., 2019; Mohtashami et al., 2023; Liu et al., 2024b; Csordás et al., 2024; Glorioso et al., 2024), and machine translation (Dabre and Fujita, 2019; Milbauer et al., 2023; Xia et al., 2019; Takase and Kiyono, 2023; Ge et al., 2022). These methods often claim to achieve comparable performance with more compact models and increased computational speed, while also setting the ground for effective adaptive compute solutions (Dehghani et al., 2019; Graves, 2016b; Schuster et al., 2021).
@@
In this paper, we also discuss how Recursive Transformers can be well suited for early-exiting techniques to accelerate decoding in LLMs. The inherent recursive structure readily enables early-exiting for individual responses within a large serving batch, which is often a practical limitation of such techniques. Vanilla Transformers encounter a synchronization issue with early-exiting, where the model must forward all layers if even a single token in a batch requires full processing (exited tokens must wait for them). Several approaches attempt to exploit this idle time by computing missing KV caches for exited tokens in later layers, which are essential for subsequent sequence generation. These techniques include state propagation (Schuster et al., 2022; Elbayad et al., 2020), SkipDecode (Del Corro et al., 2023), and parallel decoding (which can be combined with Speculative Decoding) (Bae et al., 2023; Elhoushi et al., 2024; Liu et al., 2024a; Chen et al., 2024b; Tang et al., 2024). Nevertheless, the heterogeneous parameters across varying model depths still hinder the efficient progression of exited tokens to subsequent sequences. In contrast, our Recursive Transformers enable parallel computation for tokens at different depths and sequences (in a continuous depth-wise batching paradigm)—also allow for parallel computation of missing KV caches with minimal overhead during the memory-bounded decoding phase.
@@ Appendix B Parameter Sharing Strategy
Takase and Kiyono (2023) discuss three strategies for partial layer tying in Transformer models, as depicted in Figure B.1. The SEQUENE strategy is the simplest, assigning the same parameters to consecutive layers. The CYCLE strategy repeatedly stacks a single block of unique layers to achieve the desired depth. Meanwhile, the CYCLE (REV) strategy stacks the lower layers in reverse order for the remaining layers.
In the comparative analysis of SEQUENCE and CYCLE strategies (Liu et al., 2024b), CYCLE demonstrated marginally superior zero-shot performance. Although the SEQUENCE approach, which caches shared weights (the capacity of SRAM is typically sufficient to hold a single transformer block) and computes them iteratively, has the potential to mitigate the weight transfer bottleneck between SRAM and DRAM, we prioritized compatibility with early-exiting. Consequently, we specifically employed the CYCLE strategy, which enables continuous depth-wise batching and thereby maximizes the throughput of Recursive Transformers.
@@ Appendix F Expanded Results of Initialization Methods for Looped Layers
The Average initialization method exhibited notably poor performance, particularly when applied to the Gemma model. We hypothesized that this could be attributed to instability in the model’s learned distribution, potentially arising from averaging of normalization layer weights. Relatedly, several studies (Csordás et al., 2024; Shim et al., 2024; Mohtashami et al., 2023) have explored the careful design of layer normalization in parameter-shared models. To investigate this further, we experimented with three different methods for initializing normalization weights, as outlined in Figure 1(b): averaging weights (Norm-avg), selecting weights from a single layer (Norm-choice), and zero initialization (Norm-zero). The performance trend observed among these methods varied across different model architectures. However, zero initialization of normalization layers resulted in a huge performance drop in certain architectures like TinyLlama and Pythia. Conversely, we observed no big difference between averaging and single-layer selection, suggesting that any form of distillation of the normalization weights appears to be sufficient for maintaining performance.
@@ Appendix G Expanded Results of Relaxed Recursive Transformers
Appendix H Alternative Approaches to Relaxation of Parameter Sharing
To mitigate the restrictive weight-tying inherent in parameter sharing, we employed LoRA modules as discussed in §2.3, similar to prior work (Ge et al., 2022; Shim et al., 2024). However, efficiently computing multiple LoRA modules requires specialized kernels and sequential computations of the base layers and LoRA modules, incurring computational overhead. Consequently, we explored layer-specific prompts (Liu et al., 2021) as an alternative. This approach integrates prompts specific to each layer as prefix tokens, generating layer-wise key and value states for self-attention, and is significantly more amenable to parallel computation.
Table H.1 summarizes performance of the prefix tuning method. While offering computational advantages, its reliance on small, learnable prompts resulted in limited performance gains. Additionally, without leveraging the original pretrained weights, performance was significantly lower (52.1% vs. 47.6% with the Average method in 1.07B model size). Future work will explore enhancing the effectiveness of this parallel approach, as well as other strategies such as bias term-based relaxation (Ge et al., 2022).